OpenStudy (tainted):

Add. (x5 + 7x3 - 9) + (4x5 - 2x3 + 13) 5x5 + 5x3 + 4 4x5 - 5x3 - 4 4x5 + 9x3 + 22 5x5 + 9x3 - 22

OpenStudy (tainted):

New to polynomials

OpenStudy (aaronq):

does x5 mean x^5?

OpenStudy (tainted):

yes it does

OpenStudy (aaronq):

so ill do one (x^5 + 7x^3 - 9) + (4x^5 - 2x^3 + 13) (1)break brackets.. in this case it's easy just remove them but if you had a negative sign infront of the bracket you need to change signs for everything inside that bracket eg -(1-3+5) = -1+3-5 (2) collect like terms .. group terms so you can visualize which ones are able to be added up. Since you can only add variables (in this case, "x") together when they're raised to the same power. e.g. x^2 +x^4 + 3x^2 +x^8 x^2 + 3x^2 + x^4 +x^8 = 4x^2 + x^4 + x^8 (3) write in increasing order of power

OpenStudy (aaronq):

wait i forgot to do it =(x^5 + 7x^3 - 9) + (4x^5 - 2x^3 + 13) =x^5 + 7x^3 - 9 + 4x^5 - 2x^3 + 13 = 7x^3- 2x^3+ x^5+ 4x^5 + 13 = 5x^3 + 5x^5 +13

OpenStudy (anonymous):

5x5 + 5x3 + 4

OpenStudy (aaronq):

haha oh yeah i missed the x-less number thanks for correcting me

OpenStudy (anonymous):

x5 + 7x3 - 9 + 4x5 - 2x3 + 13 x5 +4x5 +7x3-2x3 -9+13 =5x5 + 5x3 + 4

OpenStudy (anonymous):

no problem :)